# Single-loop Algorithms for Stochastic Non-convex Optimization with Weakly-Convex Constraints

### v1

@@ 4 A Hinge Exact Penalty Method and Theory
The key idea of the hinge-based exact penalty method is to solve the following unconstrained minimization [42, 13]:
$\min_{\mathbf{x}}\Phi(\mathbf{x}):=F(\mathbf{x})+\frac{\beta}{m}\sum_{k=1}^{m}[h_{k}(\mathbf{x})]_{+},$ (3)
where $[\cdot]_{+}:=\max\{0,\cdot\}$ is the hinge function, and $\beta>0$ is a penalty parameter. For clarity of notation, we define
$h_{k}^{+}(\mathbf{x})=[h_{k}(\mathbf{x})]_{+}.$
Let $\partial[h_{k}(x)]_{+}\in[0,1]$ be the subdifferential of $[z]_{+}$ at $z=h_{k}(x)$ (different from $\partial h_{k}^{+}(\mathbf{x})$).
@@
$\text{dist}\left(0,\frac{1}{m}\sum_{k=1}^{m}\partial h_{k}^{+}(\mathbf{x}^{\prime})\right)\geq\delta,\forall\mathbf{x}^{\prime}\in\mathcal{V}$ (4)
@@
$0\in\partial F(\bar{\mathbf{x}})+\frac{\beta}{m}\sum_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}})+(\bar{\mathbf{x}}-{\mathbf{x}})/\theta.$
@@
$\text{dist}\left(0,\partial F(\bar{\mathbf{x}})+\frac{\beta}{m}\sum_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}})\right)\leq\|(\bar{\mathbf{x}}-{\mathbf{x}})/\theta\|\leq\epsilon.$
Since $\partial h_{k}^{+}(\bar{\mathbf{x}})=\xi_{k}\partial h_{k}(\bar{\mathbf{x}})$ (see [2, Corollary 16.72]), where
$\xi_{k}=\begin{cases}1&\text{if }h_{k}(\bar{\mathbf{x}})>0,\\ [0,1]&\text{if }h_{k}(\bar{\mathbf{x}})=0,\\ 0&\text{if }h_{k}(\bar{\mathbf{x}})<0,\end{cases}\quad\in\partial[h_{k}(\bar{\mathbf{x}})]_{+},$
there exists $\lambda_{k}\in\frac{\beta\xi_{k}}{m}\geq 0,\forall k$ such that
$\text{dist}\left(0,\partial F(\bar{\mathbf{x}})+\sum_{k=1}^{m}\lambda_{k}\partial h_{k}(\bar{\mathbf{x}})\right)\leq\epsilon.$
Thus, we prove condition (i) in Definition 3.3. Next, let us prove condition (ii). We argue that $\max_{k}h_{k}(\bar{\mathbf{x}})\leq 0$. Suppose this does not hold, i.e., $\max_{k}h_{k}(\bar{\mathbf{x}})>0$, we will derive a contradiction. Since $\exists\mathbf{v}\in\partial F(\bar{\mathbf{x}})$ such that
$\epsilon\geq\text{dist}\left(0,\mathbf{v}+\frac{\beta}{m}\sum_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}})\right)\geq\text{dist}\left(0,\frac{\beta}{m}\sum_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}})\right)-\|\mathbf{v}\|\geq\beta\delta-L_{F},$ (5)
@@
A key assumption in Theorem 4.2 is the condition in (4). Next, we provide some sufficient conditions. For simplicity of exposition and comparison with prior works, we consider differentiable constraint functions.
@@
If there exists $\mu>0,c>0$ such that (i) $h(\mathbf{x})-\min_{\mathbf{y}}h(\mathbf{y})\leq\frac{1}{2\mu}\|\nabla h(\mathbf{x})\|^{2}$ for any $\mathbf{x}$ satisfying $h(\mathbf{x})>0$; and (ii) $\min_{\mathbf{y}}h(\mathbf{y})\leq-c$, then condition (4) holds for $\delta=\sqrt{2c\mu}$.
@@
Next, we present a sufficient condition for (4) for multiple constraints. The following lemma shows that the full rank assumption of the Jacobian matrix of $H(\mathbf{x})=(h_{1}(\mathbf{x}),\ldots h_{m}(\mathbf{x}))^{\top}$ at $x$ that violate constraints is a sufficient condition.
@@
If $\nabla H(\mathbf{x})$ is full rank for $\max_{k}h_{k}(\mathbf{x})>0$, i.e., $\lambda_{min}(\nabla H(\mathbf{x}))\geq\sigma>0,\forall\mathbf{x}\text{ such that }\max_{k}h_{k}(\mathbf{x})>0$, where $\lambda_{\min}(\cdot)$ denotes the minimum eigen-value of a matrix, then the condition (4) holds for $\delta=\sigma/m$.
We refer to the above condition as Full-Rank-at-Violating-Points condition (FRVP). We note that the FRVP condition has been used as a sufficient condition in [21] for ensuring
@@
A stronger condition for equality constrained optimization with $h_{k}(\mathbf{x})=0,k=1,\ldots,m$ was imposed in [1], where they assumed that
$\|\nabla H(\mathbf{x})^{\top}H(\mathbf{x})\|\geq\delta\|H(\mathbf{x})\|,\forall\mathbf{x}.$ (7)
To ensure this, their method requires the full-rank Jacobian assumption to hold at all points, whereas our approach only requires it at constraint-violating points. In our experiments, we verified that the condition in Lemma 4.4 holds on the trajectory of the algorithm. Detailed results are provided in Table 2 in Appendix A.
@@ 4.2 Issue with squared-hinge penalty
Let us discuss why using the squared-hinge penalty function as in [21]. Using a squared-hinge penalty function gives the following objective:
$\Phi(\mathbf{x})=F(\mathbf{x})+\frac{\beta}{m}\sum_{k=1}^{m}[h_{k}(\mathbf{x})]^{2}_{+}.$ (8)
@@
$\epsilon\geq\left\|\mathbf{v}+\frac{\beta}{m}\sum_{k=1}^{m}[h_{k}(\bar{\mathbf{x}})]_{+}\nabla h_{k}(\bar{\mathbf{x}})\right\|\geq\left\|\frac{\beta}{m}\sum_{k=1}^{m}[h_{k}(\bar{\mathbf{x}})]_{+}\nabla h_{k}(\bar{\mathbf{x}})\right\|-\|\mathbf{v}\|\geq\frac{\beta\delta}{m}\|[H(\bar{\mathbf{x}})]_{+}\|-L_{F},$
where the last step follows the assumption in (6). Since $[H(\bar{\mathbf{x}})]_{+}$ could be very small, in order to derive a contradiction we need to set $\beta$ to be very large. In [21], it was set to be the order of $O(1/\epsilon)$. As a result, $\Phi(\cdot)$ will have large weak-convexity and Lipschitz parameters of the order $O(1/\epsilon)$. This will lead to a complexity of $O(1/\epsilon^{12})$ if we follow existing studies to solve the problem in (8[16] without the smoothness of $F$ and $h_{k}$, which will be discussed in next section.
@@ 4.3 Issue of solving an equivalent equality constrained optimization
Does converting the inequality constrained problem into an equality constrained problem work? In general this answer is no. One may consider the following equivalent equality constrained problem:
@@
where $s_{1},\ldots,s_{m}$ are introduced dummy variables. A solution $\mathbf{x}^{*}$ is a KKT solution to above if there exists $\boldsymbol{\lambda}=(\lambda_{1},\ldots,\lambda_{m})^{\top}\in\mathbb{R}^{m}$ such that $0\in\partial F(\mathbf{x}^{*})+\sum_{k=1}^{m}\lambda_{k}\partial h_{k}(\mathbf{x}^{*}),$ $\lambda_{k}s_{k}=0,\forall k$, and $h_{k}(\mathbf{x}^{*})+s_{k}^{2}=0$ for $\forall{k}=1,\ldots,m$. However, this does not provide a guarantee that $\mathbf{x}_{*}$ is a KKT condition to the original constrained optimization problem as $\boldsymbol{\lambda}$ is not necessarily non-negative. For instance, consider $\min x\text{ s.t. }-1\leq x\leq 1$. After adding slack variables, we have $\min x\text{ s.t. }x+s_{1}^{2}=1,-x+s_{2}^{2}=1$. However, $(x,s_{1},s_{2})=(1,0,\sqrt{2})$ becomes a stationary point with multipliers $\lambda=(-1,0)$, but it is not a stationary point in the original problem.
While several studies [15, 11] have employed stronger second-order conditions to ensure that a KKT solution of the equality-constrained problem (9) can be converted into one for the original inequality-constrained problem, these conditions generally do not hold for the problems we consider. Moreover, when these second-order conditions fail to be satisfied, the approach of solving the equality-constrained problem (9) becomes ineffective. Our solutions will not suffer from this issue.
@@ 5 Stochastic Algorithms
In this section, we consider stochastic algorithms for solving (3) under different settings for the objective function $F(\mathbf{x})$. The first setting is the following:

### v2

@@ 4 A Hinge Exact Penalty Method and Theory
The key idea of the hinge-based exact penalty method is to solve the following unconstrained minimization (Zangwill, 1967; Evans et al., 1973), where $\beta>0$ is a penalty parameter:
$\min_{\mathbf{x}}\Phi(\mathbf{x}):=F(\mathbf{x})+\frac{\beta}{m}\sum\nolimits_{k=1}^{m}[h_{k}(\mathbf{x})]_{+},$ (2)
For clarity of notation, we define $h_{k}^{+}(\mathbf{x})=[h_{k}(\mathbf{x})]_{+}.$ Let $\partial[h_{k}(\mathbf{x})]_{+}\in[0,1]$ be the subdifferential of $z]_{+}$ at $z=h_{k}(\mathbf{x})$, and let $\partial h_{k}^{+}(\mathbf{x})$ be the subgradient in terms of $\mathbf{x}$.
@@
$\text{dist}\left(0,\frac{1}{m}\sum\nolimits_{k=1}^{m}\partial h_{k}^{+}(\mathbf{x}^{\prime})\right)\geq\delta,\forall\mathbf{x}^{\prime}\in\mathcal{V}$ (3)
@@
$0\in\partial F(\bar{\mathbf{x}})+\frac{\beta}{m}\sum\nolimits_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}})+(\bar{\mathbf{x}}-{\mathbf{x}})/\theta.$
@@
$\text{dist}\left(0,\partial F(\bar{\mathbf{x}})+\frac{\beta}{m}\sum\nolimits_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}})\right)\leq\|(\bar{\mathbf{x}}-{\mathbf{x}})/\theta\|\leq\epsilon.$
Since $\partial h_{k}^{+}(\bar{\mathbf{x}})=\xi_{k}\partial h_{k}(\bar{\mathbf{x}})$ (see Bauschke et al. (2017, Corollary 16.72)), where
$\xi_{k}=\begin{cases}1&\text{if }h_{k}(\bar{\mathbf{x}})>0,\\ [0,1]&\text{if }h_{k}(\bar{\mathbf{x}})=0,\\ 0&\text{if }h_{k}(\bar{\mathbf{x}})<0,\end{cases}\quad\subseteq\partial[h_{k}(\bar{\mathbf{x}})]_{+},$
there exists $\lambda_{k}\in\frac{\beta\xi_{k}}{m}\geq 0,\forall k$ such that $\text{dist}\left(0,\partial F(\bar{\mathbf{x}})+\sum_{k=1}^{m}\lambda_{k}\partial h_{k}(\bar{\mathbf{x}})\right)\leq\epsilon$. Thus, we prove condition (i) in Definition 3.3. Next, let us prove condition (ii). We argue that $\max_{k}h_{k}(\bar{\mathbf{x}})\leq 0$. Suppose this does not hold, i.e., $\max_{k}h_{k}(\bar{\mathbf{x}})>0$, we will derive a contradiction. Since $\exists\mathbf{v}\in\partial F(\bar{\mathbf{x}})$ such that
$\epsilon\geq\text{dist}\left(0,\mathbf{v}+\frac{\beta}{m}\sum\nolimits_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}})\right)\geq\text{dist}\left(0,\frac{\beta}{m}\sum\nolimits_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}})\right)-\|\mathbf{v}\|\geq\beta\delta-L_{F},$ (4)
@@
Assumption (3) is not specific to our method but is shared by all penalty-based approaches, including recent works such as Alacaoglu and Wright (2024), Li et al. (2024b), and Li et al. (2024a). This assumption reflects a fundamental challenge in non-convex constrained optimization: without a condition such as (3), the problem becomes ill-posed and is generally unsolvable using first-order methods. There is potential to slightly relax assumption (3). In our current formulation, we assume (3) holds over the entire set $\mathcal{V}$. However, it may suffice to require (3) to hold only on a subset of $\mathcal{V}$, defined as $\mathcal{V}_{c}:=\{x:c\geq\max_{k}h_{k}(x)>0\}$ for some constant $c>0$. This is a strictly weaker assumption.
Under this relaxation, the algorithm can be initialized within the feasible region and proceed with step sizes chosen carefully to ensure that all intermediate iterates remain within $\mathcal{V}_{c}$. Note that even if some iterates become slightly infeasible, as long as the step size is sufficiently small, the gradients of the constraint functions will guide the iterates back toward feasibility. This way, we can still conduct the same convergence analysis just select feasible initial point $\mathbf{x}_{0}$ and proceed with step sizes chosen carefully. This strategy has been successfully adopted in prior work (e.g., (Huang and Lin, 2023)).
@@
Next, we provide some sufficient conditions for (3).
@@
If there exists $\mu>0,c>0$ such that (i) $h(\mathbf{x})-\min_{\mathbf{y}}h(\mathbf{y})\leq\frac{1}{2\mu}\text{dist}(0,\partial h(x))^{2}$ for any $\mathbf{x}$ satisfying $h(\mathbf{x})>0$; and (ii) $\min_{\mathbf{y}}h(\mathbf{y})\leq-c$, then condition (3) holds for $\delta=\sqrt{2c\mu}$.
@@
Next, we present a sufficient condition for (3) for multiple constraints. The following lemma shows that the full rank assumption of the Jacobian matrix of $H(\mathbf{x})=(h_{1}(\mathbf{x}),\ldots h_{m}(\mathbf{x}))^{\top}$ at $\mathbf{x}$ that violate constraints is a sufficient condition.
@@
If $\partial H(\mathbf{x})$ is full rank for $\max_{k}h_{k}(\mathbf{x})>0$, i.e., $\text{dist}(0,\lambda_{\min}(\partial H(\mathbf{x})))\geq\sigma>0,\forall\mathbf{x}\text{ such that }\max_{k}h_{k}(\mathbf{x})>0$, where $\partial H(\mathbf{x})$ denotes the set of sub-Jacobian matrices and $\lambda_{\min}(\cdot)$ denotes the set of their minimum singular values, then the condition (3) holds for $\delta=\sigma/m$.
We refer to the above condition as Full-Rank-at-Violating-Points condition (FRVP). We note that the FRVP condition for differentiable constraint functions has been used as a sufficient condition in Li et al. (2024a) for ensuring
@@
A stronger condition for equality constrained optimization with $h_{k}(\mathbf{x})=0,k=1,\ldots,m$ was imposed in Alacaoglu and Wright (2024), where they assumed that $\|\nabla H(\mathbf{x})^{\top}H(\mathbf{x})\|\geq\delta\|H(\mathbf{x})\|,\forall\mathbf{x}$. To ensure this, their method requires the full-rank Jacobian assumption to hold at all points, whereas our approach only requires it at constraint-violating points. In our experiments, we verified that the condition in Lemma 4.4 holds on the trajectory of the algorithm. Detailed results are provided in Table 2 in Appendix A.
@@ 4.2 Issue with squared-hinge penalty
Let us discuss why using the squared-hinge penalty function as in Li et al. (2024a). For simplicity of exposition and comparison with prior works, we consider differentiable constraint functions here. Using a squared-hinge penalty function gives the following objective:
$\Phi(\mathbf{x})=F(\mathbf{x})+\frac{\beta}{m}\sum\nolimits_{k=1}^{m}[h_{k}(\mathbf{x})]^{2}_{+}.$ (6)
@@
$\epsilon\geq\left\|\mathbf{v}+\frac{\beta}{m}\sum\nolimits_{k=1}^{m}[h_{k}(\bar{\mathbf{x}})]_{+}\nabla h_{k}(\bar{\mathbf{x}})\right\|\geq\left\|\frac{\beta}{m}\sum_{k=1}^{m}[h_{k}(\bar{\mathbf{x}})]_{+}\nabla h_{k}(\bar{\mathbf{x}})\right\|-\|\mathbf{v}\|\geq\frac{\beta\delta}{m}\|[H(\bar{\mathbf{x}})]_{+}\|-L_{F},$
where the last step follows the assumption in (5). Since $H(\bar{\mathbf{x}})]_{+}$ could be very small, in order to derive a contradiction we need to set $\beta$ to be very large. In Li et al. (2024a), it was set to be the order of $O(1/\epsilon)$. As a result, $\Phi(\cdot)$ will have large weak-convexity and Lipschitz parameters of the order $O(1/\epsilon)$. This will lead to a complexity of $O(1/\epsilon^{12})$ if we follow existing studies(Hu et al., 2024) to solve the problem in (6)  without the smoothness of $F$ and $h_{k}$, which will be discussed in Section 5.
@@ 4.3 Issue of solving an equivalent equality constrained optimization
One may consider the following equivalent equality constrained problem:
@@
where $s_{1},\ldots,s_{m}$ are introduced dummy variables. A solution $\mathbf{x}^{*}$ is a KKT solution to above if there exists $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{m})^{\top}\in\mathbb{R}^{m}$ such that $0\in\partial F(\mathbf{x}^{*})+\sum_{k=1}^{m}\lambda_{k}\partial h_{k}(\mathbf{x}^{*}),$ $\lambda_{k}s_{k}=0,\forall k$, and $h_{k}(\mathbf{x}^{*})+s_{k}^{2}=0$ for $\forall{k}=1,\ldots,m$. However, this does not provide a guarantee that $\mathbf{x}_{*}$ is a KKT condition to the original constrained optimization problem as $\bm{\lambda}$ is not necessarily non-negative. For instance, consider $\min x\text{ s.t. }-1\leq x\leq 1$. After adding slack variables, we have $\min x\text{ s.t. }x+s_{1}^{2}=1,-x+s_{2}^{2}=1$. However, $(x,s_{1},s_{2})=(1,0,\sqrt{2})$ becomes a stationary point with multipliers $\lambda=(-1,0)$, but it is not a stationary point in the original problem.
While several studies (Fukuda and Fukushima, 2017; Ding and Wright, 2023) have employed stronger second-order conditions to ensure that a KKT solution of the equality-constrained problem (7) can be converted into one for the original inequality-constrained problem, these conditions generally do not hold for the problems we consider. Moreover, when these second-order conditions fail to be satisfied, the approach of solving the equality-constrained problem (7) becomes ineffective. Our solutions will not suffer from this issue.
Specifically, Proposition 3.6. in Fukuda and Fukushima (2017) shows that if a stationary solution to the equality-constrained problem (formulated using squared slack variables) satisfies the second-order sufficient condition (SOSC), then it is also a stationary solution to the original inequality-constrained problem (after simply removing the slack variables). Theorem 3.3. in reference (Ding and Wright, 2023) proves a similar result but under the weaker second-order necessary condition (SONC). To the best of our knowledge, no other conditions have been established in the literature.
Both SOSC and SONC conditions above are quite strong, as they require all first-order stationary points of the equality-constrained problem to be also second-order stationary points at the same time. Unfortunately, neither SOSC nor SONC is implied by the assumptions made in our paper. This is true even if we further assume our problem is second-order differentiable. Since our focus is on first-order methods, our assumption is related to first-order conditions only. In fact, in general, our problems may not even have gradients (non-smooth), let alone the Hessian matrix in SOSC and SONC. While the results in Fukuda and Fukushima (2017); Ding and Wright (2023) may be useful in practice as potential certificates of stationarity, their applicability is limited in practice. For example, one could first find a stationary point for the equality-constrained problem, and then verify whether SOSC or SONC holds at the obtained solution. If lucky enough and they hold, one can conclude that the solution is stationary for the original inequality-constrained problem. However, if unlucky and these second-order conditions do not hold, we do not know if the solution is stationary for the original problem. In contrast, our method directly finds a stationary point for the inequality-constrained problem without requiring such verification, hence there is no need to implement the verification procedure.
Another way is to use non-negative slack variables as following:
$\min_{\mathbf{x},s_{1},\ldots,s_{m}\geq 0}F(\mathbf{x}),~~~\text{s.t. }~~h_{k}(\mathbf{x})+s_{k}=0,k=1,\ldots,m$ (8)
where $s_{1},\ldots,s_{m}$ are introduced dummy variables. Li et al. (2021) presents some discussion how to convert an $\epsilon$-KKT solution to the above equality constrained into an $\epsilon$-KKT solution to the original inequality constrained optimization. However, such guarantee requires that boundness of the constraint functions.
@@ 5 Stochastic Algorithms
In this section, we consider stochastic algorithms for solving (2) for $F(\mathbf{x})=\mathbb{E}_{\zeta}[f(\mathbf{x},\zeta)]$, which has been considered in Ma et al. (2020); Huang and Lin (2023); Boob et al. (2023). Due to limit of space, we present extension and analysis of compositional objective $F(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}f(\mathbb{E}_{\zeta}[g_{i}(\mathbf{x},\zeta)])$ in the Appendix D. Without loss of generality, we assume the constrained functions are only accessed through stochastic samples. Below, we let $h(\mathbf{x},\mathcal{B})=\frac{1}{|\mathcal{B}|}\sum_{\xi\in\mathcal{B}}h(\mathbf{x},\xi)$ with a mini-batch of samples $\mathcal{B}$. We make the following assumptions regarding $h_{k}$, and $f$.
Assumption 5.1.
For any $\mathbf{x}$, assume there exists $h_{k}(\mathbf{x},\xi)$, and $\partial h_{k}(\mathbf{x},\xi)$, where $\xi$ is random variable, such that (a) $\mathbb{E}[h_{k}(\mathbf{x},\xi)]=h_{k}(\mathbf{x})$ and $\mathbb{E}[\partial h_{k}(\mathbf{x},\xi)]\in\partial h_{k}(\mathbf{x})$; (b) $\mathbb{E}_{\xi}[|h_{k}(\mathbf{x},\xi)-h_{k}(\mathbf{y},\xi)|^{2}]\leq L_{h}^{2}\|\mathbf{x}-\mathbf{y}\|^{2}$; (c) $\mathbb{E}_{\xi}[|h_{k}(\mathbf{x},\xi)-h_{k}(\mathbf{x})|^{2}]\leq\sigma_{h}^{2}$.
Assumption 5.2.
For any $\mathbf{x}$, assume there exists $\partial f(\mathbf{x},\zeta)$, where $\zeta$ is random variable such that $\mathbb{E}[\partial f(\mathbf{x},\zeta)]\in\partial F(\mathbf{x})$ and $\mathbb{E}_{\zeta}[\|\partial F(\mathbf{x})-\partial f(\mathbf{x},\zeta)\|^{2}]\leq\sigma_{f}^{2}$.
Algorithm 1 Algorithm for solving equation 2 under Setting I
1:Initialization: choose $\mathbf{x}_{0},\beta,\gamma_{2}$ and $\eta$.
2:for $t=0$ to T1T-1 do
3:  Sample $\mathcal{B}^{t}_{c}\subset\{1,\ldots,m\}$
4:  for each $k\in\mathcal{B}^{t}_{c}$ do
5:   Sample a mini-batch $\mathcal{B}_{2,k}^{t}$
6:   Update $u_{2,k}^{t+1}=(1-\gamma_{2})u_{2,k}^{t}+\gamma_{2}{h_{k}}(\mathbf{x}_{t};\mathcal{B}_{2,k}^{t})+\gamma_{2}^{\prime}({h_{k}}(\mathbf{x}_{t};\mathcal{B}_{2,k}^{t})-{h_{k}}(\mathbf{x}_{t-1};\mathcal{B}_{2,k}^{t}))$
7:  end for
8:  Let $u_{2,k}^{t+1}=u_{2,k}^{t},k\notin\mathcal{B}^{t}_{c}$
9:  Compute $G_{2}^{t}=\frac{\beta}{|\mathcal{B}^{t}_{c}|}\sum_{k\in\mathcal{B}^{t}_{c}}\partial h_{k}(\mathbf{x}_{t};\mathcal{B}_{2,k}^{t})[u_{2,k}^{t}]^{\prime}_{+}$, $G_{1}^{t}=\frac{1}{|\mathcal{B}^{t}|}\sum_{\zeta\in\mathcal{B}^{t}}\partial f(\mathbf{x}_{t};\zeta)$.
10:  Update $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta(G_{1}^{t}+G_{2}^{t})$
11:end for
The key to our algorithm design is to notice that the hinge-based exact penalty function $\frac{\beta}{m}\sum_{k=1}^{m}[h_{k}(\mathbf{x})]_{+}$ is a special case of non-smooth weakly-convex finite-sum coupled compositional objective as considered in Hu et al. (2024), where the inner functions are $h_{k}(\mathbf{x})$ and the outer function is the hinge function that is monotone and convex. Hence, we can directly utilize their technique for computing a stochastic gradient estimator of the penalty function.
The subgradient of the penalty function is given by $\frac{1}{m}\sum_{k=1}^{m}\beta\partial[h_{k}(\mathbf{x})]_{+}\partial h_{k}(\mathbf{x})$. However, computing $h_{k}(\mathbf{x})$ and $\partial h_{k}(\mathbf{x})$ is prohibited due to its stochastic nature or depending on many samples. To this end, we need an estimator for $h_{k}(\mathbf{x})$ and $\partial h_{k}(\mathbf{x})$. $\partial h_{k}(\mathbf{x})$ can be simply estimated by its mini-batch estimator. The challenge is to estimate $h_{k}(\mathbf{x})$ using mini-batch samples. The naive approach that estimates $h_{k}(\mathbf{x})$ by its mini-batch estimator does not yield a convergence guarantee for compositional optimization. To address this issue, we use the variance reduction technique MSVR for estimating constraint functions $h_{k}(\mathbf{x}),\forall k$ (Jiang et al., 2022), which maintains and updates an estimator $u_{2,k}$ for each constraint function $h_{k}$. We present the key steps in Algorithm 1. In particular, at the $t$-th iteration, we construct a random mini-batch $\mathcal{B}^{t}_{c}\subset\{1,\ldots,m\}$, then we update the estimators of $h_{k}(\mathbf{x}_{t}),k\in\mathcal{B}^{t}_{c}$ by Step 6 where $\gamma_{2}\in(0,1),\gamma_{2}^{\prime}=\frac{m-|\mathcal{B}_{c}|}{|\mathcal{B}_{c}|(1-\gamma_{2})}+1-\gamma_{2}$, and $\mathcal{B}^{t}_{2,k}$ is a mini-batch of samples for estimating $h_{k}(\mathbf{x}_{t})$. Then the gradient of the penalty function and the objective function at $\mathbf{x}_{t}$ can be estimated by $G_{2}^{t},G_{1}^{t}$ in Step 10, where $u_{2,k}^{t}]^{\prime}_{+}$ is the derivative of the hinge function at $u_{2,k}^{t}$. Then, we can update $\mathbf{x}_{t+1}$ by using SGD with $G_{1}^{t}+G_{2}^{t}$ as the stochastic gradient estimator.
Since the estimator $u_{2,k}$ produced by MSVR is used to approximate the true constraint value during optimization, it is useful to clarify how it behaves over iterations. Our analysis of Lemma B.3 shows that as long as $\sqrt{\gamma_{2}}\!\to\!0$, $\eta/\sqrt{\gamma_{2}}\!\to\!0$, and $t\!\to\!\infty$, the error $\lvert u_{2,k}^{t}-h_{k}(\mathbf{x}_{t})\rvert$ converges to zero. This guarantees that the solution converges to a KKT point asymptotically. For the non-asymptotic case with $\epsilon>0$, although $u_{2,k}^{t}$ may fluctuate around zero near the feasibility boundary, such oscillations can be effectively controlled by choosing sufficiently small $\eta$ and $\gamma_{2}$. This ensures that the hinge-based constraint $h_{k}(\mathbf{x}_{t})\leq\epsilon$ is reliably enforced. In our experiments, we did not observe any instability in constraint satisfaction.
The convergence of the algorithm for finding an $\epsilon$-stationary solution $\Phi_{\theta}(x)$ is given below. The proof of the theorem is presented in Appendix C.
Theorem 5.3.
Suppose Assumptions 3.1 5.1 and 5.2 hold. Let $\gamma_{2}=\mathcal{O}(\frac{|\mathcal{B}_{2,k}|\epsilon^{4}}{\beta^{4}})$, $\eta=\mathcal{O}(\frac{|\mathcal{B}_{c}||\mathcal{B}_{2,k}|^{1/2}\epsilon^{4}}{\beta^{5}m})$, after $T=\mathcal{O}(\frac{\beta^{6}m}{|\mathcal{B}_{c}||\mathcal{B}_{2,k}|^{1/2}\epsilon^{6}})$ iterations, the Algorithm 1 satisfies $\mathbb{E}[\|\nabla\Phi_{\theta}(\mathbf{x}_{\hat{t}})\|]\leq\epsilon$ from some $\theta=O(1/(\rho_{0}+\rho_{1}\beta))$, where $\hat{t}$ selected uniformly at random from $\{1,\ldots,T\}$.
Remark: While we state the convergence results for a given $\epsilon$, it is also possible to state the convergence for a given $T$. In this case, we can let $\eta,\gamma_{2}$ depend on $T$, which can yield the same complexity result. In practice, these hyper-parameters are always tuned to achieve the best performance. For example, in the first experiment we choose the initial learning rate from $\{10^{-3},10^{-4}\}$ and apply a standard decay schedule; in the second experiment we use the same magnitude choices as in Li et al. (2024a). These settings are fully reported in the experimental section, and our results show that they work well in practice.
Combining the results from Theorem 5.3 and 4.2, we demonstrate Algorithm 1 can find a nearly $\epsilon$-KKT solution to the original problem (1) in the following Corollary. The proof is provided at the end of Appendix C.
Corollary 5.4.